# Joint Beamforming Design and Bit Allocation in Massive MIMO with Resolution-Adaptive ADCs

### v2

@@ I-A Prior Works
Recent years have witnessed a proliferation of studies on massive MIMO with low-resolution ADCs (with 22224444 bits). It has been shown in [singh2009limits, jacobsson2017throughput] that a system using very few bits can approach the performance of a full-resolution one. Mezghani et al. [mezghani2012capacity] derived a closed-form lower bound for the capacity of a point-to-point MIMO system. More recent works have focused on beamforming designs [mezghani2009transmit, jacobsson2017quantized, ling2019performance, ma2024digital]. Furthermore, mixed-ADC systems, which simultaneously deploy one-bit and high-resolution ADCs, have been shown to perform better than fixed-resolution architectures, especially at high SNR [zhang2016mixed, zhang2017performance, pirzadeh2018spectral]. On the other hand, resolution-adaptive ADCs have been studied in [bai2013optimization, ahmed2017joint, choi2017resolution, nguyen2020energy, prasad2020optimizing, castaneda2021resolution, castaneda2021spawc, sheng2020energy] to flexibly balance the SE-EE tradeoff of low-resolution systems. For instance, it has been shown in [bai2013optimization, ahmed2017joint, choi2017resolution, nguyen2020energy, sheng2020energy] that efficient bit allocation strategies can offer a higher EE compared with uniform-resolution architectures. Moreover, Castañeda et al. [castaneda2021resolution] developed a resolution-adaptive fully digital receiver within an application-specific integrated circuit (ASIC). Furthermore, they demonstrated that a $256$-antenna base station with resolution-adaptive ADCs serving $16$ users allows a reduction in power consumption by $6.7$ times compared with a traditional fixed-resolution design [castaneda2021spawc]. Additional SE gains can be achieved by jointly optimizing the transmit power and ADC resolutions [prasad2020optimizing].
@@
where $\alpha$ is a constant depending on the quantizers and on the distribution of $X$, and $\eta$ represents the quantization distortion (QD). Both (1) and (2) can be employed to analyze the worst-case system performance [diggavi2001worst, hassibi2003much] assuming that $q$ or $\eta$ is a Gaussian random variable uncorrelated with $X$. Model (2) was first derived in [fletcher2007robust] and named AQNM later in [orhan2015low]; it was also called the pseudo-quantization noise model in [zhang2016mixed]. Although (2) and the Bussgang decomposition were developed from separate technical lineages, it was shown in [demir2020bussgang] that the former is the latter tailored for the case of quantization. Therefore, we call the model in (2) as the Bussgang-based AQNM (BAQNM) while we refer to (1) as the AQNM for distinction. The AQNM is typically less accurate than the BAQNM because the assumption that $q$ is uncorrelated with $X$ is generally not satisfied. In contrast, $\eta$ is uncorrelated with $X$ based on the properties of the Bussgang decomposition. Furthermore, the QD covariance matrix is a key ingredient for the performance analysis and optimization with the BAQNM. A diagonal approximation of the QD covariance matrix was derived in [mezghani2012capacity, bai2013optimization], which has since then been widely used in the literature. Nonetheless, this approximation can introduce substantial error in the performance analysis and optimization in some scenarios [demir2020bussgang, prasad2020optimizing]. It has been numerically demonstrated that, at low SNR, using the BAQNM and the approximation of the QD covariance matrix can obtain a SE close to the channel capacity [orhan2015low, roth2017achievable]. However, at high SNR, the diagonal approximation of the QD covariance matrix results in non-negligible performance overestimation, especially for massive MIMO systems with very few ADC bits [roth2017achievable, demir2020bussgang].
@@ I-B Contributions
Recently, Castañeda et. al [castaneda2021spawc, castaneda2021resolution] implemented the first fully digital systems with resolution-adaptive ADCs, where the resolution bits are dynamically adjusted to adapt to the instantaneous communications scenario, e.g., CSI and modulation scheme. They demonstrated that the adoption of resolution-adaptive ADCs can achieve power savings with several orders of magnitude, for realistic mmWave channels, underscoring the significant potential of this technology. Previous research on resolution-adaptive ADCs has mainly focused on bit allocation [bai2013optimization, choi2017resolution, nguyen2020energy] or the combined optimization of bit allocation with either transmit or receive beamforming design [ahmed2017joint, prasad2020optimizing, sheng2020energy]. However, the joint optimization of transmit-receive beamforming and bit allocation holds significant potential for achieving higher SE and providing valuable insights into the SE-EE tradeoff, yet it remains largely unexplored. Moreover, existing designs [bai2013optimization, choi2017resolution, nguyen2020energy, ahmed2017joint, prasad2020optimizing, sheng2020energy] fail to adequately address the complex coupling between transmit-receive beamformers and the bit allocation vector, which demands a comprehensive and integrated design approach. This work bridges this critical gap by introducing a joint transmit-receive beamforming design and bit allocation in point-to-point MIMO systems employing resolution-adaptive ADCs. To facilitate the design, we analyze the BAQNM and the approximation of the QD covariance matrix. Table I outlines the distinctions between our work and prior studies. The specific contributions of this paper are summarized as follows.
@@
  • We first present the fundamental properties of optimal quantization. Leveraging these properties and the Bussgang decomposition, we obtain a more accurate characterization of the BAQNM and the diagonal approximation of the QD covariance matrix, offering a new perspective compared to [mezghani2012capacity, bai2013optimization]. Our analysis shows that the BAQNM and the diagonal approximation of the QD covariance matrix only hold when Gaussian signals are optimally quantized. Furthermore, we reveal the connections between applying BAQNM and the arcsine law [jacovitti1994estimation] to one-bit quantization.
@@
  • We perform extensive numerical simulations based on the simulated QD covariance matrix, which has non-zero off-diagonal entries, using the proposed properties of optimal quantization. The results confirm the superiority of the proposed schemes over the state-of-the-art algorithms, especially beamforming alone with 22224444 ADC bits per RF chain. For example, in a $64\times 64$ MIMO system, the proposed design offers $6\%$ improvements in both SE and EE while requiring $40\%$ fewer active ADC bits compared with uniform bit allocation. Moreover, the SE-EE comparison shows that receiving more data streams with low-resolution ADCs can achieve higher SE and EE than receiving fewer data streams with high-resolution ADCs.
TABLE I: Comparison of prior works on resolution-adaptive ADCs.
Bit allocation
[bai2013optimization, choi2017resolution, nguyen2020energy]
Transmit beamforming [prasad2020optimizing]
Receive beamforming [ahmed2017joint, sheng2020energy]
Transmit-receive beamforming This work
@@ I-C Organization and Notations
Scalars, vectors, and matrices are denoted by the lowercase, boldface lowercase, and boldface uppercase letters, respectively. Furthermore, we use $\left(\cdot\right)^{*}$, $\left(\cdot\right)^{\scriptscriptstyle\mathsf{T}}$, $\left(\cdot\right)^{\scriptscriptstyle\mathsf{H}}$, and $\left(\cdot\right)^{-1}$ to represent the conjugate, transpose, conjugate transpose, and matrix inverse operators, respectively. $\left\|\cdot\right\|_{\mathcal{F}}$ signifies the Frobenius norm for matrices. In addition, the expectation and trace operators are represented by ${\mathbb{E}}\left(\cdot\right)$ and ${\rm Tr}\left(\cdot\right)$. We use $\left|a\right|$ and $\det\left(\mathbf{A}\right)$ to denote the absolute value of the scalar $a$ and the determinant of the matrix $\mathbf{A}$, respectively. The real and imaginary part operators are denoted by ${}\Re\{\cdot\}$ and ${}\Im\{\cdot\}$, respectively. Moreover, ${\rm diag}(\mathbf{a})$ or ${\rm diag}(\mathbf{A})$ returns a diagonal matrix whose diagonal entries are the same as the elements of $\mathbf{a}$ or the diagonal entries of $\mathbf{A}$. In addition, $\mathbf{A}(:,1:J)$ represents the matrix consisting of the left $J$ columns of $\mathbf{A}$. Finally, we use $\mathbf{C}_{xy}$ and $\mathbf{C}_{x}$ to represent the cross-covariance matrix between $\mathbf{x}$ and $\mathbf{y}$ and the auto-covariance matrix of $\mathbf{x}$, respectively.
@@ II-A System Model
To characterize the system performance upper bound, we assume the availability of perfect CSI at both the Rx and Tx [mo2015capacity, ling2019performance, lin2019transceiver, ning2023beamforming, dai2022delay, qi2022hybrid, zhang2017hybridly, gao2021wideband]. We note that the channel can be assumed quasi-static in some point-to-point scenarios (e.g., wireless backhaul). In addition, channel estimation with adaptive-resolution ADCs was studied in [wang2022channel]. Furthermore, an ASIC receiver integrating both the resolution-adaptive ADCs and the channel estimation module was developed in [castaneda2021resolution]. The effect of imperfect CSI will be numerically investigated in Section LABEL:sec:simulation.
II-B Signal Model with Quantization
@@
where $\mathbf{U}\in{\mathbb{C}}^{N_{\rm r}\times N_{\rm t}}$ denotes the combining matrix. Although $\mathbf{s}$ is Gaussian distributed, $\mathbf{e}$ does not follow a Gaussian distribution because of the nonlinear quantization distortion. However, we can treat the effective noise vector $\mathbf{e}$ as a Gaussian random variable and obtain a lower bound of the SE as [hassibi2003much]
@@ III-A Properties of Optimal Quantizers
Let $X$ be a real-valued, zero-mean, and unit-variance random variable, and let $Y=\sigma_{\rm y}X$. Then, we have
@@
For a zero-mean complex random variable $X=\Re\{X\}+j\Im\{X\}$ with variance $\sigma_{X}^{2}$, assume that $\Re\{X\}$ and $\Im\{X\}$ are independent and identically distributed (i.i.d.) with the same variance $\frac{\sigma_{X}^{2}}{2}$ and are independently quantized by two identical Lloyd-Max quantizers $Q(\cdot)$. With $\chi=Q(X)-X$, we obtain

### v3

@@ I-A Prior Works
Recent years have witnessed a proliferation of studies on massive MIMO with low-resolution ADCs (with 22224444 bits). It has been shown in [singh2009limits, jacobsson2017throughput] that a system using very few bits can approach the performance of a full-resolution one. Mezghani et al. [mezghani2012capacity] derived a closed-form lower bound for the capacity of a point-to-point MIMO system. More recent works have focused on beamforming designs [mezghani2009transmit, jacobsson2017quantized, ling2019performance, ma2024digital]. Furthermore, mixed-ADC systems, which simultaneously deploy one-bit and high-resolution ADCs, have been shown to perform better than fixed-resolution architectures, especially at high SNR [zhang2016mixed, zhang2017performance, pirzadeh2018spectral]. On the other hand, resolution-adaptive ADCs have been studied in [bai2013optimization, ahmed2017joint, choi2017resolution, nguyen2020energy, prasad2020optimizing, castaneda2021resolution, castaneda2021spawc, sheng2020energy, verenzuela2021optimal, verenzuela2017per] to flexibly balance the SE-EE tradeoff of low-resolution systems. For instance, it has been shown in [bai2013optimization, ahmed2017joint, choi2017resolution, nguyen2020energy, sheng2020energy, verenzuela2021optimal, verenzuela2017per] that efficient bit allocation strategies can offer a higher EE compared with uniform-resolution architectures. Moreover, Castañeda et al. [castaneda2021resolution] developed a resolution-adaptive fully digital receiver within an application-specific integrated circuit (ASIC). Furthermore, they demonstrated that a $256$-antenna base station with resolution-adaptive ADCs serving $16$ users allows a reduction in power consumption by $6.7$ times compared with a traditional fixed-resolution design [castaneda2021spawc]. Additional SE gains can be achieved by jointly optimizing the transmit power and ADC resolutions [prasad2020optimizing].
@@
where $\alpha$ is a constant depending on the quantizers and on the distribution of $X$, and $\eta$ represents the quantization distortion (QD). Both (1) and (2) can be employed to analyze the worst-case system’s performance [diggavi2001worst, hassibi2003much] assuming that $q$ or $\eta$ is a Gaussian random variable uncorrelated with $X$. Model (2) was first derived in [fletcher2007robust] and named AQNM later in [orhan2015low]; it was also called the pseudo-quantization noise model in [zhang2016mixed]. Although (2) and the Bussgang decomposition were developed from separate technical lineages, it was shown in [demir2020bussgang] that the former is the latter tailored for the case of quantization. Therefore, we call the model in (2) as the Bussgang-based AQNM (BAQNM) while we refer to (1) as the AQNM for distinction. The AQNM is typically less accurate than the BAQNM because the assumption that $q$ is uncorrelated with $X$ is generally not satisfied. In contrast, $\eta$ is uncorrelated with $X$ based on the properties of the Bussgang decomposition. Furthermore, the QD covariance matrix is a key ingredient for the performance analysis and optimization with the BAQNM. A diagonal approximation of the QD covariance matrix was derived in [mezghani2012capacity, bai2013optimization], which has since then been widely used in the literature. Nonetheless, this approximation can introduce substantial error in the performance analysis and optimization in some scenarios [demir2020bussgang, prasad2020optimizing]. It has been numerically demonstrated that, at low SNR, using the BAQNM and the approximation of the QD covariance matrix can obtain a SE close to the channel capacity [orhan2015low, roth2017achievable]. However, at high SNR, the diagonal approximation of the QD covariance matrix results in non-negligible performance overestimation, especially for massive MIMO systems with very few ADC bits [roth2017achievable, demir2020bussgang].
@@ I-B Contributions
Recently, Castañeda et. al [castaneda2021spawc, castaneda2021resolution] implemented the first fully digital systems with resolution-adaptive ADCs, where the resolution bits are dynamically adjusted to adapt to the instantaneous communications scenario, e.g., CSI and modulation scheme. They demonstrated that the adoption of resolution-adaptive ADCs can achieve power savings with several orders of magnitude for realistic mmWave channels, underscoring the significant potential of this technology. Previous research on resolution-adaptive ADCs has mainly focused on bit allocation [bai2013optimization, choi2017resolution, nguyen2020energy, verenzuela2021optimal, verenzuela2017per] or the combined optimization of bit allocation with either transmit or receive beamforming design [ahmed2017joint, prasad2020optimizing, sheng2020energy]. However, the joint optimization of transmit-receive beamforming and bit allocation holds significant potential for achieving higher SE and providing valuable insights into the SE-EE tradeoff, yet it remains largely unexplored. Moreover, existing designs [bai2013optimization, choi2017resolution, nguyen2020energy, ahmed2017joint, prasad2020optimizing, sheng2020energy, verenzuela2021optimal, verenzuela2017per] fail to adequately address the complex coupling between transmit-receive beamformers and the bit allocation vector, which demands a comprehensive and integrated design approach. This work bridges this critical gap by introducing a joint transmit-receive beamforming design and bit allocation in point-to-point MIMO systems employing resolution-adaptive ADCs.
To facilitate the design, it is essential to properly model the quantization process. Although the BAQNM is widely adopted in the literature, its foundational assumptions, which significantly influence its accuracy, have not been thoroughly examined. Existing studies mainly use the BAQNM as a plug-and-play tool without evaluating its applicability, which may lead to questionable conclusions. To address this issue, we aim to develop a more accurate quantization model that enables reliable and insightful analysis. Table I outlines the distinctions between our work and prior studies.
The specific contributions of this paper are summarized as follows.
TABLE I: Comparison of prior works on resolution-adaptive ADCs.
Reference Bit allocation Transmit beamforming Receive beamforming Quantization modeling analysis
[bai2013optimization, choi2017resolution, nguyen2020energy, verenzuela2021optimal, verenzuela2017per]
[prasad2020optimizing]
[ahmed2017joint, sheng2020energy]
This work
  • We first present the fundamental properties of optimal quantization, including the scaling law, distortion invariance, and some important statistical properties between random variable and its quantized output. The scaling law enables efficient derivation of the optimally quantized outputs for Gaussian signals with different variances, significantly reducing computational and time complexities. As a result, it allows to efficiently obtain the actual QD covariance matrix via numerical methods, which includes the non-zero off-diagonal entries. This leads to more reliable performance evaluation compared to that solely relying on the theoretical diagonal approximation. In addition, the distortion invariance property ensures that the Bussgang gain matrix in the BAQNM depends solely on the distortion factor, which is determined by the quantizer resolution. We propose a more accurate expression of this distortion factor, enabling reliable performance evaluation when employing the BAQNM.
@@
  • Leveraging the fundamental properties of optimal quantization and the Bussgang decomposition, we obtain a more accurate characterization of the BAQNM and the diagonal approximation of the QD covariance matrix compared to those developed in [mezghani2012capacity, bai2013optimization]. Our analysis shows that the BAQNM and the diagonal approximation of the QD covariance matrix only hold when Gaussian signals are optimally quantized. Moreover, we reveal the connections and nuances between applying BAQNM and the arcsine law [jacovitti1994estimation] to one-bit quantization. The BAQNM applies to only the optimal quantization, whereas the arcsine law implies more general quantization for one-bit systems.
@@
  • We perform extensive numerical simulations based on the simulated QD covariance matrix, which has non-zero off-diagonal entries, using the scaling law of optimal quantization. The results confirm the superiority of the proposed schemes over the state-of-the-art algorithms, especially beamforming alone with 22224444 ADC bits per RF chain. For example, in a $64\times 64$ MIMO system, the proposed design offers $6\%$ improvements in both SE and EE while requiring $40\%$ fewer active ADC bits compared with uniform bit allocation. Moreover, the SE-EE comparison shows that receiving more data streams with low-resolution ADCs can achieve higher SE and EE than receiving fewer data streams with high-resolution ADCs.
@@ I-C Organization and Notations
Scalars, vectors, and matrices are denoted by the lowercase, boldface lowercase, and boldface uppercase letters, respectively. Furthermore, we use $\left(\cdot\right)^{*}$, $\left(\cdot\right)^{\scriptscriptstyle\mathsf{T}}$, $\left(\cdot\right)^{\scriptscriptstyle\mathsf{H}}$, and $\left(\cdot\right)^{-1}$ to represent the conjugate, transpose, conjugate transpose, and matrix inverse operators, respectively. $\left\|\cdot\right\|_{\mathcal{F}}$ signifies the Frobenius norm for matrices, whereas $\otimes$ denotes the Kronecker product. In addition, the expectation and trace operators are represented by ${\mathbb{E}}\left(\cdot\right)$ and ${\rm Tr}\left(\cdot\right)$. We use $\left|a\right|$ and $\det\left(\mathbf{A}\right)$ to denote the absolute value of the scalar $a$ and the determinant of the matrix $\mathbf{A}$, respectively. The real and imaginary part operators are denoted by ${}\Re\{\cdot\}$ and ${}\Im\{\cdot\}$, respectively. Moreover, ${\rm diag}(\mathbf{a})$ or ${\rm diag}(\mathbf{A})$ returns a diagonal matrix whose diagonal entries are the same as the elements of $\mathbf{a}$ or the diagonal entries of $\mathbf{A}$. In addition, $\mathbf{A}(:,1:J)$ represents the matrix consisting of the left $J$ columns of $\mathbf{A}$. Lastly, we use $\mathbf{C}_{xy}$ and $\mathbf{C}_{x}$ to represent the cross-covariance matrix between $\mathbf{x}$ and $\mathbf{y}$ and the auto-covariance matrix of $\mathbf{x}$, respectively.
@@ II-A System Model
To characterize the upper bound on the system’s performance, we assume the availability of perfect CSI at both the Rx and Tx [bai2013optimization, choi2017resolution, nguyen2020energy, ahmed2017joint, prasad2020optimizing, sheng2020energy, mo2015capacity, ling2019performance, lin2019transceiver, ning2023beamforming, dai2022delay, qi2022hybrid, zhang2017hybridly, gao2021wideband]. The effect of imperfect CSI will be numerically investigated in Section LABEL:sec:imperfect_implementation. We note that the channel can be assumed quasi-static in some point-to-point scenarios (e.g., wireless backhaul). In addition, channel estimation with adaptive-resolution ADCs was studied in [wang2022channel]. Furthermore, an ASIC receiver integrating both the resolution-adaptive ADCs and the channel estimation module was implemented in [castaneda2021resolution].
II-B Channel Model
The mmWave propagation environment can be well characterized by the Saleh-Valenzuel channel model [rangan2014millimeter]. Assuming a uniform planar array (UPA) at both the Tx and the Rx, the channel matrix is expressed as [yu2016alternating, lin2019transceiver, qi2022hybrid, zhang2017hybridly]
$\mathbf{H}=\sqrt{\frac{N_{\rm t}N_{\rm r}}{N_{\rm cl}N_{\rm ray}}}\sum\limits_{i=1}^{N_{\rm cl}}\sum\limits_{l=1}^{N_{\rm ray}}\alpha_{il}\mathbf{a}_{\rm r}\left(\theta_{il}^{\rm r},\phi_{il}^{\rm r}\right)\mathbf{a}_{\rm t}^{\scriptscriptstyle\mathsf{H}}\left(\theta_{il}^{\rm t},\phi_{il}^{\rm t}\right),$ (4)
where $N_{\rm cl}$ and $N_{\rm ray}$ indicate the number of clusters and distinct rays within each cluster, respectively, and $\alpha_{il}$ denotes the gain of $l$-th ray in the $i$-th propagation cluster. In addition, $\mathbf{a}_{\rm r}\left(\theta_{il}^{\rm r},\phi_{il}^{\rm r}\right)$ and $\mathbf{a}_{\rm t}\left(\theta_{il}^{\rm t},\phi_{il}^{\rm t}\right)$ represent the receive and transmit array response vectors, respectively, where $\theta_{il}^{r}$ ($\phi_{il}^{\rm r}$) and $\theta_{il}^{\scriptscriptstyle\mathsf{T}}$ ($\phi_{il}^{\rm t}$) stand for the azimuth (elevation) angles of arrival/departure (AoAs/AoDs) of the $l$-th ray in the $i$-th propagation cluster, respectively. Assume that the Rx deploys a UPA of size $N_{\rm r,h}\times N_{\rm r,v}$ with $N_{\rm r}=N_{\rm r,h}N_{\rm r,v}$. Defining $\rho_{il}^{\rm r}=\sin(\theta_{il}^{\rm r})\sin(\phi_{il}^{\rm r})$ and $\varrho_{il}^{\rm r}=\cos(\theta_{il}^{\rm r})$, the array response vector $\mathbf{a}_{\rm r}\left(\theta_{il}^{\rm r},\phi_{il}^{\rm r}\right)$ at the Rx can be expressed as[yu2016alternating, lin2019transceiver]
$\mathbf{a}_{\rm r}\left(\rho_{il}^{\rm r},\varrho_{il}^{\rm r}\right)=\mathbf{a}_{\rm r,h}\left(\rho_{il}^{\rm r}\right)\otimes\mathbf{a}_{\rm r,v}\left(\varrho_{il}^{\rm r}\right),$ (5)
with
$\mathbf{a}_{\rm r,h}\left(\rho_{il}^{\rm r}\right)=\frac{1}{\sqrt{N_{\rm r,h}}}[1,e^{j\pi\rho_{il}^{\rm r}},\cdots,e^{j(N_{\rm r,h}-1)\pi\rho_{il}^{\rm r}}]^{\scriptscriptstyle\mathsf{T}},$ (6)
$\mathbf{a}_{\rm r,v}\left(\varrho_{il}^{\rm r}\right)=\frac{1}{\sqrt{N_{\rm r,v}}}[1,e^{j\pi\varrho_{il}^{\rm r}},\cdots,e^{j(N_{\rm r,v}-1)\pi\varrho_{il}^{\rm r}}]^{\scriptscriptstyle\mathsf{T}}.$
The array response vector $\mathbf{a}_{\rm t}\left(\theta_{il}^{\rm t},\phi_{il}^{\rm t}\right)$ at the Tx can be modeled similarly.
II-C Signal Model with Quantization
@@
where $\mathbf{U}\in{\mathbb{C}}^{N_{\rm r}\times N_{\rm t}}$ denotes the combining matrix. Although $\mathbf{s}$ is Gaussian distributed, $\mathbf{e}$ does not follow a Gaussian distribution because of the nonlinear QD. However, we can treat the effective noise vector $\mathbf{e}$ as a Gaussian random variable and obtain a lower bound of the SE as [hassibi2003much]
@@ III-A Properties of Optimal Quantizers
Let $X$ be a real-valued, zero-mean, and unit-variance Gaussian random variable, and let $Y=\sigma_{\rm y}X$. Then, we have
@@
For a zero-mean complex random variable $X=\Re\{X\}+j\Im\{X\}$ with variance $\sigma_{X}^{2}$, assume that $\Re\{X\}$ and $\Im\{X\}$ are independent and identically distributed (i.i.d.) with the same variance $\frac{\sigma_{X}^{2}}{2}$ and are independently quantized with two identical Lloyd-Max quantizers $Q(\cdot)$. With $\chi=Q(X)-X$, we obtain